# Exploring Exponential Runge-Kutta Methods: A Survey

### v2

@@ 4. Recent advances
The year 2011 marks a decisive milestone in the study of exponential methods, thanks to the publication of the already mentioned paper by Al-Mohy and Higham [al2011computing]. As explained in Sect. 2.2, this work provided a fundamental improvement in the field by introducing innovative algorithms for computing the action of the matrix exponential on a vector without explicitly calculating the entire matrix exponential, improving both computational efficiency and numerical stability.
@@ 4.1. Exponential methods of high stiff order
The first part of this section is devoted to summarize a series of papers published by Vu Thai Luan and various co-authors. They provide specific order conditions, fundamental to construct the numerical methods. We recall here their most important contributions, and refer back to each paper for detailed derivations and calculations.
@@
In 2014, Vu Thai Luan and Alexander Ostermann coauthored three papers. In [luan2014exponential], they considered expRosenbrock methods under the reformulation in Hochbruck et al. [hochbruck2009exponential]. The authors addressed the convergence analysis in a different way: namely, instead of placing the exact solution into the numerical scheme and working with defects, they analyzed the local error first. This new approach provided a more systematic procedure to derive expRosenbrock methods of higher stiff order, which the authors used to construct families of methods of orders 3, 4 and 5 (see [luan2014exponential, Table 1]). In [luan2014stiff], the authors applied the approach developed for expRosenbrock methods in [luan2014exponential] to build order conditions up to 5 (whereas the literature at the time provided them only up to order 4) for ExpRK methods (see [luan2014stiff, Table 1]). In [luan2014explicit], the authors proved that, in the case of constant step size, one can reach stiff order 5 by only satisfying the order 5 condition in the weak form, as it was proved in [hochbruck2005explicit] for orders up to 4. They concluded by building a method with stiff order 5, which needs 8 stages. We refer to the recapitulatory table [luan2014explicit, Table 1]. In 2016, the same authors [luan2016parallel] built more accurate and efficient expRosenbrock of (stiff) order 4 and 5 (compared to their collaboration in 2014), and of order 6 (see [luan2016parallel, Table 1]). The method exploits parallelism by minimizing the number of sequential stages, meaning that if two of the stages can be computed independently, then they count as one, greatly improving the corresponding computational time.
@@
The following year, Luan [luan2017fourth] independently built a 2-stage order 4 expRosenbrock for time-dependent PDEs. It is the second superconvergent (here, meaning $p=2s$) exponential method ever built, after expRosenbrock-Euler. The performance was tested against other expRosenbrock methods and MATLAB’s ode15s on systems which are either non-stiff or semilinear with non-stiff nonlinear part.
@@
In 2020, Luan [luan2021efficient] proposed a construction for more accurate and efficient ExpRK methods of (stiff) order 4 and 5, compared to the existing ones obtained in [hochbruck2005explicit] and [luan2014explicit]. This was achieved by minimizing the number of sequential stages, similarly to what was done in [luan2016parallel] for expRosenbrock. Moreover, he exploited both results from [al2011computing]: efficient algorithm for computing $\mathrm{e}^{tA}b$ and augmented matrix $\tilde{A}$ such that $\mathrm{e}^{\tilde{A}}$ is equivalent to computing linear combinations of $\varphi$-functions. To do so, the new methods are built such that the $i$-th stage is defined by a single linear combination of $\varphi_{k}(c_{i}hA)$. The resulting ExpRK of order 4 has 6 stages instead of 5, but 4 sequential stages. The 5th order one has now 10 stages instead of 8, but 5 sequential stages. Refer to [luan2021efficient, Table 1 & Sect. 6] for a detailed explanation. The implementation of the methods makes use of their own code phipm_simul_iom, publicly available. In the same year, Luan et al. [luan2020new] introduced multirate ExpRK, using different time scales for the linear term (fast) and the nonlinear term (slow) of a semilinear (possibly abstract) ODE. The methods need to be such that the $i$-th stage is defined by a single linear combination of terms of the form $\varphi_{k}(c_{i}hA)$ (as in [luan2021efficient]) which is a nontrivial requirement. The idea is to approximate the nonlinear term in both internal and final stages with a polynomial built assuming a small timestep $h$ and then solve the ODE with the polynomials replacing the nonlinear terms, using a large timestep $H$. The idea builds on [hochbruck2011adams], where this approach was applied to exponential multistep methods (of Adams type).
@@
Shifting our focus to the most recent advances, in 2024, Luan et al. [luan2024efficient] explored applications to gene regulatory networks, which are defined by a certain number of semilinear ODEs where the linear part is diagonal and the nonlinear part can be more or less stiff. In the examples considered in the paper, the ODEs are low dimensional, the linear part may be stiff in the sense of highly oscillatory and the nonlinear one is only moderately stiff (in these models, the nonlinear term can become very stiff when many gene products are involved). The authors build an order 2 exponential method which is Rosenbrock-like, in the sense that derivatives of the right-hand side appear in the coefficients. They compare accuracy (error vs. $h$) and efficiency (error vs. CPU time) with other exponential methods and classical RK. The proposed method is the most efficient and second most accurate after expRosenbrock-Euler. The better performance than classical RK may be partly due to the matrix being diagonal (thus, exponential being even easier to compute). This latter study thus aligns with the literature on the application of ExpRK methods to mathematical models in biology and physics. A detailed analysis of the relevant contributions will be provided in the final part of this section. In the same year, Luan and Alhsmy [luan2024sixth] developed the first order 6 ExpRK methods for the time integration of stiff parabolic PDEs, leveraging the exponential B-series theory extended by Luan and Ostermann for stiff problems [luan2013exponential]. They derived 36 stiff order conditions, see [luan2024sixth, Table 1], and remarked that despite these methods requiring a large number of stages, their computational cost is similar to that of a 6 stages method.
@@ 4.2. Exponential methods for DDEs
However, the presence of the delay term may affect stability of the underlying numerical methods [belzen03], which has motivated researches to investigate the stability of ExpRK methods for DDEs of the form (9). Xu, Zhan and Sui [xu2011stability] study the case of a scalar, linear equation of the form
@@
and analyze stability properties of EERK methods when applied to DDEs with a discrete delay, for which (10) plays the role of a test function, and its delayed term is treated as nonlinear. Whenever $|\lambda|+\mathrm{Re}(\mu)<0$, the solution of (10) is asymptotically stable [inthout1992] and $P$-stable methods for DDEs are those which preserve this property when $h$ is such that $\tau=mh$ for an integer $m>0$. A numerical method for DDEs is $GP$-stable if it preserves asymptotic stability regardless of the value of $h<\tau$. As shown in [xu2011stability], the concepts of $P$- and $GP$-stability coincide in the case of ExpRK methods for DDEs. Some years later, Zhao, Zhan and Ostermann [zhao2016stability] studied various contractivity and stability concepts for EERK methods in the context of semilinear DDEs of the form
@@
Whenever $|\lambda|+\mathrm{Re}(\mu)\leq 0$, the solution of (10) is contractive, that is, it satisfies $|u(t)|\leq\|\phi\|_{\infty}$ for all $t\geq t_{0}$. $P$-contractive methods for DDEs are those which preserve this property when $h$ is such that $\tau=mh$ [torelli1989contractivity] and $GP$-contractive methods for DDEs are those which preserve this property for any choice of $h$. Zhan and Ostermann provide sufficient conditions for $GP$-contractivity of EERK methods and prove $GP$-contractivity for various methods of order 1 and 2. They also extend the concept of $GP$-stability to nonautonomous equations, which is satisfied by the Magnus method described in [gonzalez2006magnus]. Finally, they analyze $RN$- and $GRN$-stability, introduced again in [torelli1989contractivity], which also sufficient conditions on $L$, $g$ in (11) such that, for all u,vu,\,v solutions of (11) with initial conditions $u_{t_{0}}=\phi,\,v_{t_{0}}=\psi$, the following holds:
@@
Zhao, Zhan and Xu [zhao2018d] analyzed $D$-convergence. An ExpRK method for DDEs is $D$-convergent of order $p$ if, when applied to (11) with initial condition $y_{0}=\phi$, it satisfies
@@
Besides stability, some research has been devoted to the accuracy of ExpRK methods for DDEs. In particular, Fang and Zhan [fang2021high] follow the abstract approach in [hochbruck2005explicit] to derive stiff order conditions up to order 5 for semilinear DDEs of the form (11), and prove the analogous convergence theorem and support the results with numerical tests on PDEs whose semi-discretization in space gives DDEs of the target form.
More recently, Andò and Vermiglio [ando2024exponential, ando2024proceedings] investigated the application of ExpRK methods for the time-integration of general DDEs of the form (8), as well as renewal equations (REs), namely, equations of the form
@@
They consider, in particular, the reformulation of DDEs and REs as abstract (ordinary) differential equations (ADEs), which comes from the sun-star theory [dgg07, DiekmannBook]. The ADEs obtained are semilinear regardless of the right-hand sides of the original delay equations, which makes them amenable of time-integration via ExpRK methods. Since the abstract setting provided by the sun-star calculus includes both DDEs and REs, the goal in [ando2024exponential] is to define a unifying framework to study the derivation of ExpRK methods for delay equations, as well as their convergence properties. Another approach to solve problems defined by DDEs is to discretize the resulting ADEs into a finite-dimensional system of ODEs and apply ExpRK methods to the latter, as done in [ando2024proceedings]. Therein, the ODEs are obtained via pseudospectral discretization [bdgsv16], which gives rise to a stiff linear term, with eigenvalues having large negative real parts. Several numerical tests with EXPINT, a MATLAB package for exponential integrators [expint07] confirm the validity of the method.
The most recent contribution on the applications of ExpRK methods to DDEs is represented by the extension of the ExpRK methods of collocation type introduced in [hochbruck2005exponential] to semilinear (abstract) parabolic problems such that the nonlinear term is defined by a discrete delay [huang2025timedependent]. The delay can be time-dependent, as long as it is non-vanishing and that the map $t\mapsto\tau(t)$ is strictly increasing. This produces a mesh of non-equispaced breaking points, i.e., points of discontinuity of some derivative of the solution. The authors propose two ways to extend the methods in [hochbruck2005exponential] to the new classes of equations, both taking the breaking points into account to define a continuous approximation of the solution in a piecewise fashion. They prove that the methods are convergent with order $s$, the number of stages, and obtain superconvergence results provided some additional restrictions on the parameters are satisfied.
@@ 4.3. Exponential methods for mathematical modeling
In this third and final part, we present an analysis of the literature employing ExpRK methods in modeling contexts, which has garnered significant interest in recent years. [^9] For the reader’s convenience, given that the equations pertain to physical models, we have opted to retain the original notation used in each cited paper throughout this section. The first contribution we examine is from 2011 and is authored by Dimarco and Pareschi [dimarco2011exponential]. In this work, the authors, motivated by the computational challenges arising from rarefied gas dynamics, introduces a class of ExpRK methods for stiff kinetic equations, with particular emphasis on the homogeneous Boltzmann equation. In the fluid regime, the stiffness of the problem makes standard explicit methods inefficient, as the equation takes the form
@@
with $\lambda=\mu\Delta t/\varepsilon$ and suitable coefficients $W_{i}(\lambda)$. This actually ensures unconditional stability, correct asymptotic preservation, and key properties such as positivity and entropy conservation, making these methods suitable for both deterministic schemes and Monte Carlo simulations. Finally, the authors highlight how their approach is closely related to the IF methods, sharing the same mathematical structure but offering greater flexibility in designing high-order schemes. In 2014, Michels, Sobottka and Weber [michels2014exponential] applied ExpRK methods to the modeling framework of Newton’s equations of motion for elastodynamic systems. These systems are described by second-order differential equations of the form
@@
It is in the last two years, however, that we have witnessed a significant increase in the number of studies on the application of ExpRK methods across various domains. This growing body of research reflects the increasing interest in these numerical techniques and their potential advantages. In the following, we will provide a comprehensive analysis of the current state of the art in this promising field of numerical analysis. In 2024, Caliari, Cassini, Einkemmer and Ostermann [Cassini23] proposed an fundamental acceleration technique for ExpRK integrators and Lawson schemes to efficiently solve semilinear advection-diffusion-reaction equations. The model under consideration is of the form:
@@
Again in 2024, multiple works were published on the subject of order reduction of ExpRK methods, focusing specifically on problems with non-commutative operators. In a first work, Hoang [hoang2024order] studied the application of ExpRK methods to nonlinear parabolic equations with non-commutative operators. The model under consideration is an equation of the form:
@@
In a second second work authored by Dang and Hoang [dang2024avoid], the authors investigated the issue of order reduction in third-order ExpRK methods applied to problems with non-commutative operators. In particular, they proved that if specific order conditions are not satisfied, the method’s accuracy may drop to approximately $2.5$ instead of $3$. To overcome this problem, the authors propose a 4 stages ExpRK method that meets all the necessary conditions to preserve the expected order. The studied method takes the form:
@@
with the coefficients $a_{ij}(-\tau A)$ and $b_{i}(-\tau A)$ defined through linear combinations of the functions $\varphi_{k}(-\tau A)$. A detailed convergence analysis is conducted, based on a semigroup formulation in a Banach space, showing that adopting higher regularity on the initial data helps to avoid order reduction. Numerical tests confirm the validity of the approach, demonstrating significant improvements in accuracy and stability compared to standard three-stage methods. [^11] In an additional work, Einkemmer, Hoang, and Ostermann [einkemmer2024should] study the efficiency of exponential integrators for advection-dominated problems. They compare Leja and Krylov-based methods for computing matrix exponential actions, evaluating their performance against explicit RK schemes. Their results show that while exponential integrators perform similarly to explicit methods in fully advection-dominated regimes, they offer advantages when parts of the domain exhibit diffusion. Among the tested approaches, Leja-based methods prove more efficient than Krylov iterations, particularly on modern supercomputers where inner product computations are costly.
In 2024, Wang, Zhao, and Liao [wang2024exponential] presented a unified theoretical framework for analyzing the energy dissipation properties of three classes of RK methods applied to Phase Field Crystal (PFC) models, which describe crystalline dynamics at atomic scales. The fundamental model is given by the gradient flow:

### v3

@@ 4. Recent advances
The year 2011 marks a decisive milestone in the study of exponential methods, thanks to the publication of the already mentioned paper by Al-Mohy and Higham [2]. As explained in Sect. 2.2, this work provided a fundamental improvement in the field by introducing innovative algorithms for computing the action of the matrix exponential on a vector without explicitly calculating the entire matrix exponential, improving both computational efficiency and numerical stability.
@@ 4.1. Exponential methods of high stiff order
The first part of this section is mostly devoted to summarize a series of papers published by Vu Thai Luan and various co-authors. They provide specific order conditions, fundamental to construct the numerical methods. We recall here their most important contributions, and refer back to each paper for detailed derivations and calculations.
@@
In 2014, Vu Thai Luan and Alexander Ostermann coauthored three papers. In [70], they considered expRosenbrock methods under the reformulation in Hochbruck et al. [48]. The authors addressed the convergence analysis in a different way: namely, instead of placing the exact solution into the numerical scheme and working with defects, they analyzed the local error first. This new approach provided a more systematic procedure to derive expRosenbrock methods of higher stiff order, which the authors used to construct families of methods of orders 3, 4 and 5 (see [70, Table 1]). In [71], the authors applied the approach developed for expRosenbrock methods in [70] to build order conditions up to 5 (whereas the literature at the time provided them only up to order 4) for ExpRK methods (see [71, Table 1]). In [69], the authors proved that, in the case of constant step size, one can reach stiff order 5 by only satisfying the order 5 condition in the weak form, as it was proved in [49] for orders up to 4. They concluded by building a method with stiff order 5, which needs 8 stages. We refer to the recapitulatory table [69, Table 1]. In 2016, the same authors [72] built more accurate and efficient expRosenbrock of (stiff) order 4 and 5 (compared to their collaboration in 2014), and of order 6 (see [72, Table 1]). The method exploits parallelism by minimizing the number of sequential stages, meaning that if two of the stages can be computed independently, then they count as one, greatly improving the corresponding computational time.
@@
The following year, Luan [74] independently built a 2-stage order 4 expRosenbrock for time-dependent PDEs. It is the second superconvergent (here, meaning $p=2s$) exponential method ever built, after expRosenbrock-Euler. The performance was tested against other expRosenbrock methods and MATLAB’s ode15s on systems which are either non-stiff or semilinear with non-stiff nonlinear part.
@@
In 2020, Luan [75] proposed a construction for more accurate and efficient ExpRK methods of (stiff) order 4 and 5, compared to the existing ones obtained in [49] and [69]. This was achieved by minimizing the number of sequential stages, similarly to what was done in [72] for expRosenbrock. Moreover, he exploited both results from [2]: efficient algorithm for computing $\mathrm{e}^{tA}b$ and augmented matrix $\tilde{A}$ such that $\mathrm{e}^{\tilde{A}}$ is equivalent to computing linear combinations of $\varphi$-functions. To do so, the new methods are built such that the $i$-th stage is defined by a single linear combination of terms $\varphi_{k}(c_{i}hA)$, where $\{c_{i}\}_{i}$ are the abscissas of the method. The resulting ExpRK of order 4 has 6 stages instead of 5, but 4 sequential stages. The 5th order one has now 10 stages instead of 8, but 5 sequential stages. Refer to [75, Table 1 & Sect. 6] for a detailed explanation. The implementation of the methods makes use of their own code phipm_simul_iom, publicly available. In the same year, Luan et al. [67] introduced multirate ExpRK, using different time scales for the linear term (fast) and the nonlinear term (slow) of a semilinear (possibly abstract) ODE. The methods need to be such that the $i$-th stage is defined by a single linear combination of terms of the form $\varphi_{k}(c_{i}hA)$ (as in [75]) which is a nontrivial requirement. The idea is to approximate the nonlinear term in both internal and final stages with a polynomial built assuming a small time step $h$ and then solve the ODE with the polynomials replacing the nonlinear terms, using a large time step $H$. The idea builds on [52], where this approach was applied to exponential multistep methods (of Adams type).
@@
Shifting our focus to the most recent advances, in 2024, Luan et al. [73] explored applications to gene regulatory networks, which are defined by a certain number of semilinear ODEs where the linear part is diagonal and the nonlinear part can be more or less stiff. In the examples considered in the paper, the ODEs are low dimensional, the linear part may be stiff in the sense of highly oscillatory and the nonlinear one is only moderately stiff (in these models, the nonlinear term can become very stiff when many gene products are involved). The authors build an order 2 exponential method which is Rosenbrock-like, in the sense that derivatives of the right-hand side appear in the coefficients. They compare accuracy (error vs. $h$) and efficiency (error vs. CPU time) with other exponential methods and classical RK. The proposed method is the most efficient and second most accurate after expRosenbrock-Euler. The better performance than classical RK may be partly due to the matrix being diagonal (thus, exponential being even easier to compute). This latter study thus aligns with the literature on the application of ExpRK methods to mathematical models in biology and physics. A detailed analysis of the relevant contributions will be provided in the final part of this section. In the same year, Luan and Alhsmy [66] developed the first order 6 ExpRK methods for the time integration of stiff parabolic PDEs, leveraging the exponential B-series theory extended by Luan and Ostermann for stiff problems [68]. They derived 36 stiff order conditions, see [66, Table 1], and remarked that despite these methods requiring a large number of stages, their computational cost is similar to that of a 6 stages method.
We conclude this section with a very recent development concerning (high-order) Lawson schemes. In these schemes, internal stages typically require matrix-vector multiplications with matrices of the form $\mathrm{e}^{(c_{j}-c_{i})hA}$, where the $\{c_{i}\}_{i}$ are again the abscissas of the method. For this reason, having equispaced abscissas is a desirable property, since it requires the computation of only one matrix exponential. Lawson observed this in his seminal paper [63] with a focus on RK4, while Golden [37] recently explored the idea in detail. Using the Newton-Raphson method to establish the coefficients, he arrived at developing a sixth order method with eight stages. Finally, he provided evidence of its efficiency by testing it on the 2D incompressible Navier-Stokes equation.
@@ 4.2. Exponential methods for DDEs
However, the presence of the delay term may affect stability of the underlying numerical methods [5], which has motivated researches to investigate the stability of ExpRK methods for DDEs of the form (9). Xu, Zhan and Sui [94] study the case of a scalar, linear equation of the form
@@
and analyze stability properties of EERK methods when applied to DDEs with a discrete delay, for which (10) plays the role of a test function, and its delayed term is treated as nonlinear. Whenever $|\lambda|+\mathrm{Re}(\mu)<0$, the solution of (10) is asymptotically stable [54] and $P$-stable methods for DDEs are those which preserve this property when $h$ is such that $\tau=mh$ for an integer $m>0$. A numerical method for DDEs is $GP$-stable if it preserves asymptotic stability regardless of the value of $h<\tau$. As shown in [94], the concepts of $P$- and $GP$-stability coincide in the case of ExpRK methods for DDEs. Some years later, Zhao, Zhan and Ostermann [97] studied various contractivity and stability concepts for EERK methods in the context of semilinear DDEs of the form
@@
Whenever $|\lambda|+\mathrm{Re}(\mu)\leq 0$, the solution of (10) is contractive, that is, it satisfies $|u(t)|\leq\|\phi\|_{\infty}$ for all $t\geq t_{0}$. $P$-contractive methods for DDEs are those which preserve this property when $h$ is such that $\tau=mh$ [91] and $GP$-contractive methods for DDEs are those which preserve this property for any choice of $h$. Zhan and Ostermann provide sufficient conditions for $GP$-contractivity of EERK methods and prove $GP$-contractivity for various methods of order 1 and 2. They also extend the concept of $GP$-stability to nonautonomous equations, which is satisfied by the Magnus method described in [38]. Finally, they analyze $RN$- and $GRN$-stability, introduced again in [91], which also sufficient conditions on $L$, $g$ in (11) such that, for all u,vu,\,v solutions of (11) with initial conditions $u_{t_{0}}=\phi,\,v_{t_{0}}=\psi$, the following holds:
@@
Zhao, Zhan and Xu [98] analyzed $D$-convergence. An ExpRK method for DDEs is $D$-convergent of order $p$ if, when applied to (11) with initial condition $y_{0}=\phi$, it satisfies
@@
Besides stability, some research has been devoted to the accuracy of ExpRK methods for DDEs. In particular, Fang and Zhan [34] follow the abstract approach in [49] to derive stiff order conditions up to order 5 for semilinear DDEs of the form (11), and prove the analogous convergence theorem and support the results with numerical tests on PDEs whose semi-discretization in space gives DDEs of the target form.
More recently, Andò and Vermiglio [4, 3] investigated the application of ExpRK methods for the time-integration of general DDEs of the form (8), as well as renewal equations (REs), namely, equations of the form
@@
They consider, in particular, the reformulation of DDEs and REs as abstract (ordinary) differential equations (ADEs), which comes from the sun-star theory [27, 28]. The ADEs obtained are semilinear regardless of the right-hand sides of the original delay equations, which makes them amenable of time-integration via ExpRK methods. Since the abstract setting provided by the sun-star calculus includes both DDEs and REs, the goal in [4] is to define a unifying framework to study the derivation of ExpRK methods for delay equations, as well as their convergence properties. Another approach to solve problems defined by DDEs is to discretize the resulting ADEs into a finite-dimensional system of ODEs and apply ExpRK methods to the latter, as done in [3]. Therein, the ODEs are obtained via pseudospectral discretization [11], which gives rise to a stiff linear term, with eigenvalues having large negative real parts. Several numerical tests with EXPINT, a MATLAB package for exponential integrators [8] confirm the validity of the method.
The most recent contribution on the applications of ExpRK methods to DDEs is represented by the extension of the ExpRK methods of collocation type introduced in [50] to semilinear (abstract) parabolic problems such that the nonlinear term is defined by a discrete delay [53]. The delay can be time-dependent, as long as it is non-vanishing and that the map $t\mapsto\tau(t)$ is strictly increasing. This produces a mesh of non-equispaced breaking points, i.e., points of discontinuity of some derivative of the solution. The authors propose two ways to extend the methods in [50] to the new classes of equations, both taking the breaking points into account to define a continuous approximation of the solution in a piecewise fashion. They prove that the methods are convergent with order $s$, the number of stages, and obtain superconvergence results provided some additional restrictions on the parameters are satisfied.
@@ 4.3. Exponential methods for mathematical modeling
In this third and final part, we present an analysis of the literature employing ExpRK methods in modeling contexts, which has garnered significant interest in recent years. [^9] For the reader’s convenience, given that the equations pertain to physical models, we have opted to retain the original notation used in each cited paper throughout this section. The first contribution we examine is from 2011 and is authored by Dimarco and Pareschi [29]. In this work, the authors, motivated by the computational challenges arising from rarefied gas dynamics, introduce a class of ExpRK methods for stiff kinetic equations, with particular emphasis on the homogeneous Boltzmann equation. In the fluid regime, the stiffness of the problem makes standard explicit methods inefficient, as the equation takes the form
@@
with $\lambda=\mu\Delta t/\varepsilon$ and suitable coefficients $W_{i}(\lambda)$. This actually ensures unconditional stability, correct asymptotic preservation, and key properties such as positivity and entropy conservation, making these methods suitable for both deterministic schemes and Monte Carlo simulations. Finally, the authors highlight how their approach is closely related to the IF methods, sharing the same mathematical structure but offering greater flexibility in designing high-order schemes. In 2014, Michels, Sobottka and Weber [78] applied ExpRK methods to the modeling framework of Newton’s equations of motion for elastodynamic systems. These systems are described by second-order differential equations of the form
@@
It is in the last years, however, that we have witnessed a significant increase in the number of studies on the application of ExpRK methods across various domains. This growing body of research reflects the increasing interest in these numerical techniques and their potential advantages.
Recently, renewed attention has been devoted to the numerical integration of the nonlinear Schrödinger equation, particularly in connection with high-order exponential integrators and structure-preserving techniques. In 2022, the Schrödinger equation was revisited through a scalar auxiliary variable (SAV) formulation combined with exponential Runge-Kutta integrating factor methods, allowing the construction of linearly implicit schemes of order up to four that exactly preserve a discrete modified energy functional [56]. The SAV approach is based on the introduction of an additional scalar variable associated with the nonlinear part of the Hamiltonian, which enables efficient time integration while avoiding the solution of nonlinear systems. The proposed methods were shown to be competitive in terms of accuracy versus computational cost when compared with other SAV-based and energy-oriented schemes. [^10] We note that the SAV-based schemes preserve a modified energy rather than the original Hamiltonian of the Schrödinger equation. While this modification does not affect the continuous problem and is shown numerically to keep the true energy well controlled over moderate time intervals, it may lead to qualitative differences in long-time simulations. The authors justify this choice by the resulting linear implicitness, high-order accuracy, and favorable efficiency-error balance. More recently, Li and Li [64] proposed multiple relaxation exponential Runge-Kutta methods for the nonlinear Schrödinger equation, extending the relaxation framework to exponential integrators. Their approach introduces several relaxation parameters that are determined at each time step so as to enforce the exact preservation of multiple invariants of the continuous problem, including mass and energy, while maintaining arbitrary high-order accuracy. The resulting schemes remain explicit in the linear part and avoid nonlinear solvers, while achieving simultaneous invariant preservation at the fully discrete level. In a different but related direction, Ji, Li, and Ostermann [55] developed low regularity exponential-type integrators for the derivative nonlinear Schrödinger equation. Their analysis is based on a careful reformulation of the equation that isolates resonant interactions and allows for sharp error bounds under minimal regularity assumptions on the solution. The proposed methods are shown to converge without requiring classical Sobolev smoothness, thereby addressing a major limitation of standard exponential integrators when applied to derivative nonlinearities.
Various authors have focused on the development of high-order time integration schemes for gradient flows and semilinear parabolic equations that can simultaneously achieve high accuracy and preserve intrinsic structural properties, such as energy dissipation, equilibria, and inequality constraints. ExpRK methods are particularly attractive due to their stability properties, but their practical implementation often faces challenges related to the use of exponential operators, including excessive damping of high-frequency modes and difficulties in maintaining structure-preserving features at higher order. Within this context, Teng and Zhang [90] proposed structure-preserving reconstructions of integrating factor RK schemes and developed an exponential-free framework achieving up to third-order accuracy. Wang et al. [92] further extended this line of research by constructing an up-to-fourth-order ExpRK framework for a large class of phase-field models, and by establishing unconditional energy dissipation for the resulting schemes. Their approach replaces the exponential functions in the integrating factor RK framework with recursive approximations, or with a combination of exponential and linear functions, thereby avoiding the excessive damping of high-frequency modes while preserving equilibria. These works complement earlier studies [95, 96], in which fourth-order reconstructions of integrating factor RK schemes were developed to preserve maximum principles or more general inequality constraints.
Starting from 2024, the literature has seen several contributions aimed at improving the efficiency of exponential integrators for specific yet widespread classes of problems, as well as at investigating and mitigating order reduction effects. Within this line of research, Caliari, Cassini, Einkemmer and Ostermann [14] proposed a fundamental acceleration technique for ExpRK integrators and Lawson schemes to efficiently solve semilinear advection-diffusion-reaction equations. The model under consideration is of the form:
@@
Again in 2024, multiple works were published on the subject of order reduction of ExpRK methods, focusing specifically on problems with non-commutative operators. In a first work, Hoang [45] studied the application of ExpRK methods to nonlinear parabolic equations with non-commutative operators. The model under consideration is an equation of the form:
@@
In a second work authored by Dang and Hoang [24], the authors investigated the issue of order reduction in third-order ExpRK methods applied to problems with non-commutative operators. In particular, they proved that if specific order conditions are not satisfied, the method’s accuracy may drop to approximately $2.5$ instead of $3$. To overcome this problem, the authors propose a 4 stages ExpRK method that meets all the necessary conditions to preserve the expected order. The studied method takes the form:
@@
with the coefficients $a_{ij}(-\tau A)$ and $b_{i}(-\tau A)$ defined through linear combinations of the functions $\varphi_{k}(-\tau A)$. A detailed convergence analysis is conducted, based on a semigroup formulation in a Banach space, showing that adopting higher regularity on the initial data helps to avoid order reduction. Numerical tests confirm the validity of the approach, demonstrating significant improvements in accuracy and stability compared to standard three-stage methods.
In an additional work, Einkemmer, Hoang and Ostermann [31] studied the efficiency of exponential integrators for advection-dominated problems. They compared Leja and Krylov-based methods for computing matrix exponential actions, evaluating their performance against explicit RK schemes. Their results show that while exponential integrators perform similarly to explicit methods in fully advection-dominated regimes, they offer advantages when parts of the domain exhibit diffusion. Among the tested approaches, Leja-based methods proved more efficient than Krylov iterations, particularly on modern supercomputers where inner product computations are costly. The paper focused on performance modeling: counting core operations to understand theoretical cost ratios between exponential integrators and explicit schemes. The recent complementary work [25], instead, measured real CPU time in numerical experiments. It showed Krylov-based methods can work better for small steps in real CPU usage – a nuanced result not apparent from the theoretical model.
In 2024, Wang, Zhao, and Liao [93] presented a unified theoretical framework for analyzing the energy dissipation properties of three classes of RK methods applied to phase-field Crystal (PFC) models, which describe crystalline dynamics at atomic scales. The fundamental model is given by the gradient flow: